# Multi-LF: A Unified Continuous Learning Framework for Real-Time DDoS Detection in Multi-Environment Networks

### v1

@@ III-C4 GNB
GNB Default Default -
III-D Phase 4: Multi-LF with Continuous Learning
In Multi-LF, we use two models: M1 (a lightweight model trained on a small sample of the dataset) and M2 (a complex model trained on the full dataset). Initially, data is fed into the M1 model, which predicts the traffic label. If M1’s prediction confidence is 100%, the model proceeds to the next sample for prediction or ends the process. If M1’s confidence is below 100%, the control shifts to M2 and the real-time traffic is passed to M2.
M2 then makes its prediction. If M2’s confidence is above 90%, M1 is retrained under a continuous learning mechanism using the predicted label from M2. If M2’s confidence is also below 90%, the control is handed over to a human expert for interaction. The expert monitors the specific traffic and predicts the label. Then, M1 is retrained using the new traffic data and the label provided by the human expert. Similarly, human interaction is maintained with M2 to update it periodically. This process ensures continuous monitoring and improvement of both models’ performance as shown in Figure 5.
Figure 5: Multi-LF Flow Diagram
@@ III-D Phase 4: Multi-LF with Continuous Learning
Algorithm 1 describes the continuous learning process. Here, $D$ represents the dataset, and $CT$ denotes the confidence threshold set at 0.9. $PL$ refers to the predicted label’s output by the models, while $T$ indicates the incoming real-time traffic. The algorithm employs two models: $M1$, a lightweight model trained on a small sample ($SS$) of the dataset for quick initial predictions, and $M2$, a complex model trained on the full dataset of 40.58 million instances for secondary predictions. During prediction, $S$ represents the current traffic packet. $p1$ and $c1$ are the prediction and confidence from $M1$, respectively. If $c1>CT$, the prediction ($p1$) is accepted as the output ($O$). If not, the sample is passed to $M2$ for prediction. $p2$ and $c2$ are the prediction and confidence from $M2$. If $c2>CT$, $M1$ is retrained incrementally with $S$ and $p2$. If $c2\leq CT$, a human expert provides a prediction ($p_{h}$), and both models are retrained with $S$ and $p_{h}$. This process ensures continuous model improvement and adaptation.
Algorithm 1 Continuous Learning with M1 and M2 Models
1:Input: $D$, $CT=0.9$
2:Output: $PL$ for incoming $T$
@@
3:Initialize: Train $M1$ on $SS$ of 400K instances using K-Means clustering; Train $M2$ on the full dataset of 40.58M instances
4:while $T$ exists do
5:     $S=\text{next\_sample}(T)$
6:     $p1,c1=M1.\text{predict}(S)$
7:     if $c1>CT$ then
8:         $O=p1$
@@
10:         $p2,c2=M2.\text{predict}(S)$
11:         if $c2>CT$ then
12:              $O=p2$
13:              $M1.\text{retrain}(S,p2)$
@@
15:              $p_{h}=\text{get\_human\_label}(S)$
16:              $O=p_{h}$
17:              $M1.\text{retrain}(S,p_{h})$
18:              $M2.\text{retrain}(S,p_{h})$triangleright Periodic human updates
@@
20:     end if
21:     Store or utilize $O$
22:end while
IV Results & Discussion
In our experimentation, the evaluation and training of models were conducted on a Core i7 11th generation machine with a Windows operating system. The system has 64GB of RAM and a 1 TB SSD. We use a VMWare workstation to run Ubuntu 24.04, where we run all experiments. We used the scikit-learn library for the implementation of models and NS3 version 3.36.1.
We evaluated all models in terms of accuracy, precision, recall, F1 score, computational time, and memory usage. Accuracy is defined as the proportion of correctly predicted samples to the total number of samples, calculated as follows:
$Acc.=\frac{TP+TN}{TP+TN+FP+FN}$ (16)
In the context of malicious attack detection, precision is the ratio of correctly predicted malicious traffic (true positives) to the total traffic instances predicted as malicious (true positives plus false positives). It measures how many of the detected malicious activities are actually malicious. It is calculated as follows:
$Precision=\frac{TP}{TP+FP}$ (17)
Recall, or sensitivity, in malicious attack detection, is the ratio of correctly predicted malicious traffic (true positives) to all actual malicious traffic instances (true positives plus false negatives). It measures how many of the actual malicious activities are detected by the model. It is calculated as follows:
$Recall=\frac{TP}{TP+FN}$ (18)
The F1 Score is the harmonic mean of precision and recall, providing a balance between the two metrics, especially useful in scenarios with imbalanced datasets. Malicious attack detection gives a single score that represents the balance between precision and recall. It is calculated as follows:
$F1Score=2\cdot\frac{Precision\cdot Recall}{Precision+Recall}$ (19)
IV-A Lightweight-M1 Models Results Used for Continuous Learning
Table IV presents the results for the ML models, referred to as M1, which were trained on a small sample consisting of 400,000 instances. We used the partial fit method to train four models, with and without min-max scaling. Without scaling, the performance of the models was poor, except for BNB, which achieved an accuracy score of 1.000, and MNB, which reached 0.972. In contrast, SGDC and Perceptron completely failed. This is because SGDC and Perceptron are linear models that are sensitive to the scale of features. When features have different scales, the gradient descent optimization process becomes inefficient, leading to poor convergence. The models struggle to learn effectively from unscaled data, resulting in poor performance across most classes.
TABLE IV: Results for M1 Models using M-En Data
With Scaling Model Accuracy Class Precision Recall F1-score SGDC 1.000 IoT.B 1.00 1.00 1.00 IoT.M 1.00 1.00 1.00 Tr.B 1.00 1.00 1.00 Tr.M 1.00 1.00 1.00 Macro Avg. 1.00 1.00 1.00 Perceptron 1.000 IoT.B 1.00 1.00 1.00 IoT.M 1.00 1.00 1.00 Tr.B 1.00 1.00 1.00 Tr.M 1.00 1.00 1.00 Macro Avg 1.00 1.00 1.00 MNB 1.000 IoT.B 1.00 1.00 1.00 IoT.M 1.00 1.00 1.00 Tr.B 1.00 1.00 1.00 Tr.M 1.00 1.00 1.00 Macro Avg 1.00 1.00 1.00 BNB 1.000 IoT.B 1.00 1.00 1.00 IoT.M 1.00 1.00 1.00 Tr.B 1.00 1.00 1.00 Tr.M 1.00 1.00 1.00 Macro Avg 1.00 1.00 1.00 Without Scaling Accuracy Class Precision Recall F1-score 0.250 IoT.B 0.25 1.00 0.40 IoT.M 0.00 0.00 0.00 Tr.B 0.00 0.00 0.00 Tr.M 0.00 0.00 0.00 macro avg 0.06 0.25 0.10 0.250 IoT.B 0.00 0.00 0.00 IoT.M 0.00 0.00 0.00 Tr.B 0.25 1.00 0.40 Tr.M 0.00 0.00 0.00 macro avg 0.06 0.25 0.10 0.972 IoT.B 0.99 1.00 1.00 IoT.M 1.00 1.00 1.00 Tr.B 1.00 0.90 0.95 Tr.M 0.91 0.99 0.95 macro avg 0.97 0.97 0.97 1.000 IoT.B 1.00 1.00 1.00 IoT.M 1.00 1.00 1.00 Tr.B 1.00 1.00 1.00 Tr.M 1.00 1.00 1.00 macro avg 1.00 1.00 1.00
@@ IV-A Lightweight-M1 Models Results Used for Continuous Learning
On the other hand, the performance of all models with scaling was excellent across all evaluation metrics, with all models achieving scores of 1.000. Scaling ensures that all features contribute proportionately to the model, enhancing the training efficiency and effectiveness of SGDC and Perceptron classifiers. It normalizes the feature space, resulting in more balanced weight updates and better model performance. For instance, if packet sizes range from 0 to 10,000 and another feature like traffic rate ranges from 0 to 1, the larger feature disproportionately influences the model’s parameters, leading to suboptimal performance. Scaling mitigates this problem by normalizing the features. Similarly, MNB and BNB models assume feature independence and use probabilities based on feature occurrences. Without scaling, the probability estimates can become skewed, affecting the model’s ability to correctly classify instances. After scaling, these models show improved performance due to the normalized feature space.
Table V presents the performance of various ML models, referred to as M1, trained on a 400,000-sample dataset using 10-fold cross-validation. Each model was evaluated with and without min-max scaling. Without scaling, the SGDC and Perceptron models performed poorly, achieving an accuracy of only 0.25 with a standard deviation (Std) of 0.0. This suboptimal performance is due to the models’ sensitivity to feature scale; the gradient descent optimization process becomes inefficient when features have different scales, leading to poor convergence and low precision (0.0625 $\pm$ 0.0), recall (0.25 $\pm$ 0.0), and F1 scores (0.1 $\pm$ 0.0). Without scaling, even though the Std is low, the accuracy and other evaluation metrics are also low.
In contrast, with min-max scaling, both SGDC and Perceptron achieved perfect accuracy scores of 1.0 with a Std of 0.0 across all metrics. Scaling ensures that all features contribute proportionately, normalizing the feature space and resulting in balanced weight updates and better model performance. For instance, features like packet size and traffic rate need normalization to prevent one from disproportionately influencing the model parameters. Similarly, the MNB and BNB models also showed significant improvements with scaling. Without scaling, MNB achieved a mean accuracy of 0.9719 with a Std of 0.0328, while BNB reached a mean accuracy of 0.9992 with a Std of 0.0023. After scaling, both models achieved perfect scores (1.0 $\pm$ 0.0). The results show the significance of BNB in both cases, with and without scaling, so we use it as the M1 model in our proposed approach.
@@
TABLE V: K-Fold Cross Validation Results for M1 Models using M-En Data
Model Metric Fold 1 Fold 2 Fold 3 Fold 4 Fold 5 Fold 6 Fold 7 Fold 8 Fold 9 Fold 10 Mean $\pm$ Std SGDC + Scaling Accuracy 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 Precision 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 Recall 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 F1-score 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 SGDC Accuracy 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 ± 0.0 Precision 0.0625 0.0625 0.0625 0.0625 0.0625 0.0625 0.0625 0.0625 0.0625 0.0625 0.0625 ± 0.0 Recall 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 ± 0.0 F1-score 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1 ± 0.0 Perceptron + Scaling Accuracy 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 Precision 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 Recall 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 F1-score 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 Perceptron Accuracy 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 ± 0.0 Precision 0.0625 0.0625 0.0625 0.0625 0.0625 0.0625 0.0625 0.0625 0.0625 0.0625 0.0625 ± 0.0 Recall 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25 ± 0.0 F1-score 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1 ± 0.0 MNB + Scaling Accuracy 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 Precision 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 Recall 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 F1-score 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 MNB Accuracy 0.9922 0.9311 0.9736 0.9018 0.9991 0.9995 0.9990 0.9440 0.9978 0.9815 0.9719 ± 0.0328 Precision 0.9924 0.9453 0.9759 0.9289 0.9991 0.9995 0.9990 0.9539 0.9978 0.9825 0.9774 ± 0.0246 Recall 0.9922 0.9311 0.9736 0.9018 0.9991 0.9995 0.9990 0.9440 0.9978 0.9815 0.9719 ± 0.0328 F1-score 0.9922 0.9299 0.9735 0.8979 0.9991 0.9994 0.9990 0.9433 0.9977 0.9814 0.9714 ± 0.0339 BNB + Scaling Accuracy 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 Precision 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 Recall 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 F1-score 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $\pm$ 0.0 BNB Accuracy 0.9923 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 0.9992 ± 0.0023 Precision 0.9925 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 0.9993 ± 0.0022 Recall 0.9923 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 0.9992 ± 0.0023 F1-score 0.9923 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 0.9992 ± 0.0023
IV-B Complex-M2 Models Results
@@
The performance of M2 models is shown in Table VI. These results are based on a full dataset consisting of approximately 40 million samples, and we experimented only after data scaling because, in the M1 case, we obtained the best results with data scaling. The performance of all models is significant, with RF achieving a perfect 1.000 accuracy score and outperforming all other evaluation metrics. RF manages many features and determines their importance, which is particularly beneficial in network traffic data that often contains diverse and numerous features. By focusing on the most relevant features, RF enhances predictive accuracy for identifying malicious traffic in our M-En network. Furthermore, RF scales well with large datasets, as in our case, and handles extensive data efficiently, a common requirement in network traffic analysis. While BNB performed poorly with an accuracy score of 0.8316, all models took considerable time to train but achieved significant results. Our proposed approach used RF as M2 because of its significant performance.
TABLE VI: Results for M2 Models using M-En Data
Model Accuracy Class Precision Recall F1-score
LR 0.9995 IoT.B 1.00 1.00 1.00
@@
Macro Avg 1.00 1.00 1.00
The performance metrics for the M2 models in terms of correct prediction (CP), wrong prediction (WP), and error rate are presented in Table VII. RF model demonstrated the most significant performance with the highest CP of 10,146,052 and the lowest error rate of 0.0000, which justifies its selection as M2 in our proposed approach. LR and Perceptron models both performed admirably, with error rates of 0.0005, indicating their reliability in malicious traffic detection. The GNB and ADA models also showed strong performance with error rates of 0.0016 and 0.0013, respectively. On the other hand, the BNB model had a significantly higher error rate of 0.1684, indicating its limited effectiveness in this context. The MNB model also showed a relatively higher error rate of 0.0057 compared to other models. These results highlight the robustness of the RF model and the varying degrees of effectiveness among different ML models in detecting malicious traffic.
TABLE VII: M2 Performance Metrics
@@
LR 10140676 5395 0.0005
SGDC 10128182 17889 0.0018
Perceptron 10140625 5446 0.0005
MNB 10088644 57427 0.0057
BNB 8437084 1708987 0.1684
GNB 10130116 15955 0.0016
RF 10146052 19 0.0000
ADA 10132709 13362 0.0013
Table VIII shows the computational time required for the training and testing of models used at the M1 and M2 levels, measured in seconds. The M2 models were trained on a larger dataset and thus exhibited higher computational times compared to the M1 models. In the M2 category, LR and ADA took significantly longer, with times of 13,815.6875 seconds and 14,106.234375 seconds, respectively. RF also required considerable time, 12,841.046875 seconds, reflecting its complex ensemble nature. Comparatively, models like SGDC, Perceptron, MNB, and BNB showed much lower computational times, highlighting their relative efficiency for the same tasks. However, RF is significant in terms of accuracy and also average computational cost, so we chose it for the proposed approach from M2 and BNB from M1.
TABLE VIII: M1 & M2 Computational Time
@@
IV-C Results in Real-Time Environment
In this section, we present the results of Multi-LF in real-time scenarios. For real-time testing, we used NS-3 simulation to generate sample data. The data was collected centrally and then passed to our traffic analyzer, the MTDS framework. We conducted experiments under different scenarios to evaluate the framework’s performance. We collected packets over specific time windows and then passed the batch of packets to the framework to determine the accuracy. This process was repeated over several iterations, and the average scores were reported.
We select the best models from online testing, evaluate their performance, and compare them across different scenarios. Scenario 1: M1 is used for attack detection without continuous learning capabilities and evaluated with DDoS and benign traffic. Scenario 2: M1 is used for attack detection with continuous learning capabilities. Scenario 3: M2 is used independently for testing. Scenario 4: M1 and M2 are deployed together with Multi-LF without human involvement. Scenario 5: M1 and M2 are deployed with Multi-LF and human involvement. The implementation of Scenarios 4 and 5 in Multi-LF is illustrated in Figure 6.

### v2

@@ 3.3.1 Baseline M1 & M2:
[…]
3.3.2 Proposed DL-Based M1 & M2:
In this study, we propose two DL-based architectures for M1 and M2: a lightweight Transformer (T1 $\rightarrow$ M1), a complex Transformer (T2 $\rightarrow$ M2), a lightweight LSTM (L1 $\rightarrow$ M1), and a complex LSTM (L2 $\rightarrow$ M2). We evaluate these models against several baselines. The key differences between the lightweight and complex architectures lie in the size of the training datasets and the number of layers in each architecture.
[…]
Figure 3: Proposed DL-based T2 Model Architecture
[…]
All models in this study were compiled using the Adam optimizer and the categorical_crossentropy loss function, which is suitable for multi-class classification tasks. The models were trained for 10 epochs with a batch size of 128. These settings were consistently applied across all Transformer- and LSTM-based architectures.
3.4 Phase 4: Multi-LF Framework
In the Multi-LF, we utilize two models: M1 (a lightweight model trained on a small subset of the dataset) and M2 (a more complex model trained on the full dataset). Initially, incoming data is passed to M1 for traffic label prediction. If M1’s prediction confidence is 90%, it proceeds to the next sample or terminates the process. If M1’s confidence is below 90%, the control shifts to M2, which then performs its own prediction. If M2’s confidence exceeds 90%, M1 is retrained using the predicted label from M2 through a continuous learning mechanism. To mitigate the catastrophic forgetting problem during this retraining, we employ a weight interpolation technique, where the updated weights $\mathbf{w}_{\text{new}}$ are computed as: $\mathbf{w}_{\text{new}}=\alpha\mathbf{w}_{\text{updated}}+(1-\alpha)\mathbf{w}_{\text{old}},$with $\alpha=0.75$, balancing between the new and old weights. The parameter $\alpha$ controls the trade-off between model adaptability and knowledge retention. In this work, $\alpha=0.75$ is selected to preserve 75% of prior information while allowing adaptation to new traffic patterns [rustam2025adaptive]. This value was determined empirically through a series of trials with $\alpha\in{0.25,0.50,0.75,1.0}$. Lower values (0.25 and 0.50) caused the model to overfit recent samples and forget previously learned patterns, whereas $\alpha=1.0$ completely inhibited adaptation to new data. Since our continuous learning setup receives relatively small batches of incoming traffic, $\alpha=0.75$ provided the most stable and consistent performance, and this configuration is reported in our results. Further, if M2’s confidence is also below 90%, control is transferred to a human expert, who labels the traffic sample. M1 is then retrained using the new data and the expert-provided label. This approach enables continuous learning, expert interaction, and ongoing performance enhancement of both models, as illustrated in Figure 4.
Figure 4: Overview of Multi-LF Flow Diagram
@@ 3.4 Phase 4: Multi-LF Framework
Algorithm 1 describes the continuous learning process. Let $\mathcal{T}$ denote the incoming traffic stream, and $x$ represent a current traffic sample. The models used are $M_{1}$ (a lightweight model) and $M_{2}$ (a complex model), trained on datasets $\mathcal{D}_{s}$ (a small subset) and $\mathcal{D}$ (the full dataset), respectively. During prediction, $M_{1}$ and $M_{2}$ produce predicted labels $\hat{y}_{1}$ and $\hat{y}_{2}$, with associated confidence scores $\gamma_{1}$ and $\gamma_{2}$. If confidence is low, a human expert provides a label $\hat{y}_{h}$. The final output label is denoted by $\hat{y}$, and the confidence threshold for decision-making is set as $\theta=0.9$.
Algorithm 1 Proposed Multi-LF Algorithm
1:Input: Dataset $\mathcal{D}$, Confidence Threshold $\theta=0.9$
2:Output: Predicted label $\hat{y}$ for incoming traffic stream $\mathcal{T}$
@@
3:Initialize: Train model $M_{1}$ on subset $\mathcal{D}_{s}$ (400K samples using K-Means); train model $M_{2}$ on full dataset $\mathcal{D}$ (40.58M samples)
4:while $\mathcal{T}$ is not empty do
5:  $x\leftarrow\text{next\_sample}(\mathcal{T})$
6:  $\hat{y}_{1},\gamma_{1}\leftarrow M_{1}.\text{predict}(x)$
7:  if $\gamma_{1}\leftarrow\theta$ then
8:   $\hat{y}\leftarrow\hat{y}_{1}$
@@
10:   $\hat{y}_{2},\gamma_{2}\leftarrow M_{2}.\text{predict}(x)$
11:   if $\gamma_{2}>\theta$ then
12:     $\hat{y}\leftarrow\hat{y}_{2}$
13:     $M_{1}.\text{retrain}(x,\hat{y}_{2})$
@@
15:     $\hat{y}_{h}\leftarrow\text{get\_human\_label}(x)$
16:     $\hat{y}\leftarrow\hat{y}_{h}$
17:     $M_{1}.\text{retrain}(x,\hat{y}_{h})$ $\triangleright$ Periodic human updates
18:   end if
@@
20:  Store or use $\hat{y}$
21:end while
4 Evaluation and Results Discussion
We conduct all training and evaluation on an Intel Core i7-11800H host equipped with 64 GB RAM and a 1 TB SSD. The experiments run inside an Ubuntu 24.04 virtual machine (VMware Workstation) with Python 3.12 and scikit-learn 1.6. To emulate realistic conditions, we replay mixed IoT and traditional traffic through DDoShield-IoT and classify this traffic on a promiscuous Docker node that captures the entire packet stream. We report accuracy, precision, recall, F1 score, runtime, memory footprint, and estimated CO2 emissions (via CodeCarbon [courty2024codecarbon]) for experiments.
4.1 Baseline Lightweight M1 Models Results
@@
Table 5 shows results for the baseline M1 models, trained on a 400,000-instance sample using partial fit, with and without min-max scaling. Without scaling, the performance of the models is poor, except for BNB, which achieves an accuracy score of 1.000, and MNB, which reaches an accuracy score of 0.972. In contrast, both SGDC and PER have a very low accuracy scale. These results occur since SGDC and PER are linear models that are sensitive to the scale of features. When features have different scales, the gradient descent optimization process becomes inefficient, leading to poor convergence. The models struggle to learn effectively from unscaled data, resulting in poor performance across most classes. However, the performance of all models with scaling is excellent in all evaluation metrics, with all models achieving scores of 1.00. Scaling ensures that all features contribute proportionately to the model, enhancing the training efficiency and effectiveness of SGDC and PER classifiers. Additionally, scaling normalizes the feature space, leading to more balanced weight updates and improved model performance. For instance, if packet sizes range from 0 to 10,000 and another feature, such as traffic rate, ranges from 0 to 1, the larger feature disproportionately influences the model’s parameters, leading to suboptimal performance. Here, scaling mitigates this problem by normalizing the features. Similarly, the MNB and BNB models assume feature independence and use probabilities based on the occurrence of features. Without scaling, the probability estimates can become skewed, affecting the model’s ability to correctly classify instances.
Table 5: Baseline M1 Models Results Using M-En Data
[…]
Figure 5 shows the 10-fold cross-validation performance of M1. Each model was evaluated with and without min-max scaling. Without scaling, SGDC and PER performed poorly (F1 = 0.10 $\pm$ 0.00) due to their sensitivity to feature magnitude, which hinders gradient descent convergence. After scaling, both achieved perfect F1 = 1.00 $\pm$ 0.00. Scaling normalizes feature ranges, such as packet size and traffic rate, ensuring balanced weight updates and improved performance. MNB and BNB also benefited from scaling: MNB improved from 0.97 $\pm$ 0.04 to 1.00 $\pm$ 0.00, while BNB maintained perfect performance in both cases, demonstrating its robustness. Consequently, BNB is selected as the M1 model in our approach.
@@
Figure 5: Baseline M1 Models F1 Score using K-Fold Cross Validation
To validate the effectiveness of our feature extraction configuration and threshold settings, we conducted experiments using three distinct parameter scenarios. Each scenario varied the processing interval (PI), abnormal size threshold (AST), port frequency threshold (PFT), and short-lived threshold (SLT), all of which directly influence flow-based feature computation from PCAP files. As shown in Figure 6, all models achieved perfect accuracy under our balanced configuration, confirming its suitability. PER, SGDC, and MNB exhibited strong robustness across scenarios, while BNB showed greater sensitivity to threshold changes. These results underscore the importance of carefully tuning feature aggregation parameters and validate the reliability and generalizability of our selected configuration.
Figure 6: Accuracies Across Three Feature Extraction configurations
4.2 Baseline M2 Models Results
@@
The performance of M2 models is shown in Table 6. These results are based on a full dataset consisting of approximately 40 million samples, and we experimented only after data scaling because, in the M1 case, we obtained the best results with data scaling. The performance of all models is significant, with RF achieving a perfect 1.000 accuracy score and outperforming all other evaluation metrics. RF manages many features and determines their importance, which is particularly beneficial in network traffic data that often contains diverse and numerous features. By focusing on the most relevant features, RF enhances predictive accuracy for identifying malicious traffic in our M-En network. Furthermore, RF scales well with large datasets, as in our case, and handles extensive data efficiently, a common requirement in network traffic analysis. While BNB performed poorly with an accuracy score of 0.8316, all models took considerable time to train but achieved significant results. Our baseline approach for Multi-LF  used RF as M2 because of its significant performance.
Table 6: Results For Baseline M2 Models Using M-En Data
Model Acc Class Pre Rec F1
[…]
@@
Avg 1.00 1.00 1.00
The performance metrics for the M2 models in terms of correct prediction (CP), wrong prediction (WP), and error rate are presented in Table 7. RF model demonstrated the most significant performance with the highest CP of 10,146,052 and the lowest error rate of 0.0000, which justifies its selection as M2 in our proposed approach. LR and PER models both performed admirably, with error rates of 0.0005, indicating their reliability in malicious traffic detection. The GNB and ADA models also showed strong performance with error rates of 0.0016 and 0.0013, respectively. On the other hand, the BNB model had a significantly higher error rate of 0.1684, indicating its limited effectiveness in this context. The MNB model also showed a relatively higher error rate of 0.0057 compared to other models. These results highlight the robustness of the RF model and the varying degrees of effectiveness among different ML models in detecting malicious traffic.
Table 7: Baseline M2 Models Performance Metrics
@@
[…]
SGDC 10,128,182 17,889 0.0018
[…]
MNB 10,088,644 57,427 0.0057
[…]
GNB 10,130,116 15955 0.0016
[…]
ADA 10,132,709 13362 0.0013
Table 8 shows the computational time required for the training and testing of baseline models used at the M1 and M2 levels, measured in seconds. The M2 models were trained on a larger dataset and thus exhibited higher computational times compared to the M1 models. In the M2 category, LR and ADA took significantly longer, with times of 13,815.6875 seconds and 14,106.234375 seconds, respectively. RF also required considerable time, 12,841.046875 seconds, reflecting its complex ensemble nature. Comparatively, models like SGDC, PER, MNB, and BNB showed much lower computational times, highlighting their relative efficiency for the same tasks. However, RF is significant in terms of accuracy and also average computational cost, so we chose it for the proposed approach from M2 and BNB from M1.
Table 8: Baseline M1 & M2 Models Computational Time
@@
4.3 Results in Real-Time Environment
In this section, we present the results of Multi-LF in real-time scenarios. For real-time testing, we used NS-3 simulation to generate sample data. The data was collected centrally and then passed to Multi-LF. We conducted experiments under different scenarios to evaluate the framework’s performance. We collected packets over specific time windows and then passed the batch of packets to the framework to determine the accuracy. This process was repeated over several iterations, and the average scores were reported. We select the best models from online testing, evaluate their performance, and compare them across different scenarios. (i) Scenario 1:M1 is used for attack detection without continuous learning capabilities and evaluated with benign and malicious traffic. (ii) Scenario 2:M1 is used for attack detection with continuous learning capabilities. (iii) Scenario 3:M2 is used independently for testing. (iv) Scenario 4:M1 and M2 are deployed together under Multi-LF without human involvement. (v) Scenario 5:M1 and M2 are deployed under Multi-LF with human involvement. The implementation of Scenarios 4 and 5 in Multi-LF is illustrated in Figure 7.